# Evaluating the Effect of Retrieval Augmentation on Social Biases

### v1

@@ Appendix C Additional Accuracy Evaluations
GPT-3.5 45.24 / 75.74 27.58 / 73.12 39.93 / 74.21 31.80 / 73.07
Llama3-8B-Inst. 50.30 / 60.52 24.36 / 63.89 31.99 / 65.82 28.03 / 64.53
Qwen-7B-Inst. 81.45 / 52.03 15.43 / -2.08 16.27 / -1.39 15.87 / -0.10
Qwen-14B 63.79 / 82.84 49.75 / 77.08 45.88 / 84.52 47.57 / 78.27
Qwen-14B-Inst. 96.53 / 71.92 89.78 / 67.76 92.66 / 73.12 87.85 / 71.33
Table 10: Comparison of ambiguous / disambiguous accuracy (separated by ‘/’) when using different retrieval methods to retrieving documents from the full-set. For each generator LLM, maximum and minimum accuracy are shown respectively in red and blue.
@@
Appendix D Prompt Sensitivity
We test the sensitivity of the prompt that we use to evaluate social biases using BBQ in this section. Specifically, we check the sensitivity of the prompt when using few-shot examples in the prompt.
Recent studies demonstrate that LLMs can exhibit robust few-shot performance on a variety of downstream tasks, where one or more examples are provided to guide the model in a specific text generation task. In our experiments, we randomly selected eight gender bias related instances from the English BBQ dataset as examples, including four ambiguous questions and four disambiguated questions (two with stereotyped contexts and two with counter-stereotyped contexts). We include the selected few-shot examples at the beginning of the instruction prompt shown in Figure 5.
As shown in Table 11, incorporating few-shot examples does not always reduce the Diff-Bias scores compared to the zero-shot setting, neither in the w/o RAG condition nor when top-10 documents are retrieved (i.e. w/ RAG) from the stereo-set using VectorIndex. For example, for GPT-3.5-turbo, the ambiguous Diff-Bias scores increase from 5.16 (w/o RAG, zero-shot) to 15.43 under the few-shot setting, and for the w/ RAG, from 14.53 to 25.20. Similar trends can be observed across other models.
On the other hand, as indicated in Table 12, few-shot prompting improves the accuracy, particularly for the ambiguous questions. For instance, GPT-3.5-turbo’s accuracy in ambiguous contexts w/o RAG rises from 45.24 in the zero-shot setting to 60.17 with few-shot. This suggests that although few-shot prompting can help the model to better understand the task, it does not significantly affect the social biases induced by the retrieved stereotype set or the model’s inherent social biases.
Model w/o RAG w/ RAG
Zero-shot Few-shot Zero-shot Few-shot
GPT-3.5 5.16 / -9.33 15.43 / 4.37 14.53 / 7.14 25.20 / 9.82
Llama3-8B-Inst. 5.65 / 1.59 2.83 / 0.20 14.68 / -0.40 6.30 / 1.79
Qwen-7B-Inst. 10.02 / -3.67 9.38 / 7.44 24.01 / 0.50 21.78 / 6.85
Qwen-14B 3.77 / -7.34 3.32 / 2.68 13.99 / -2.68 14.63 / 7.04
Qwen-14B-Inst. -2.38 / -2.38 -1.14 / -5.46 4.61 / 2.68 2.43 / 1.88
Table 11: Diff-Bias scores for the ambiguous and disambiguated contexts (values separated by ‘/’) under different prompting strategies. In each group (“w/o RAG” and “w/ RAG”), for ambiguous and disambiguated values separately, the diff-bias with the lowest absolute value is highlighted in bold.
Model w/o RAG w/ RAG
Zero-shot Few-shot Zero-shot Few-shot
GPT-3.5 45.24 / 75.74 60.17 / 79.32 27.83 / 71.68 26.19 / 81.15
Llama3-8B-Inst. 50.30 / 60.52 65.82 / 53.57 26.19 / 63.74 54.51 / 60.07
Qwen-7B-Inst. 81.45 / 52.03 83.09 / 48.67 62.60 / 53.72 70.98 / 49.36
Qwen-14B 63.79 / 82.84 84.08 / 76.79 46.83 / 77.63 52.73 / 81.75
Qwen-14B-Inst. 96.53 / 71.92 97.97 / 75.55 87.05 / 68.30 96.08 / 64.63
Table 12: Accuracy for the ambiguous and disambiguated contexts (values separated by ‘/’) under different prompting strategies. In each group (“w/o RAG” and “w/ RAG”), for ambiguous and disambiguated values separately, the highest value is highlighted in bold.

### v2

@@ Appendix F Additional Accuracy Evaluations
GPT-3.5 45.24±2.17 / 75.74±1.87 27.58±1.95 / 73.12±1.94 39.93±2.14 / 74.21±1.91 31.80±2.03 / 73.07±1.94
Llama3-8B-Inst. 50.30±2.18 / 60.52±2.13 24.36±1.87 / 63.89±2.10 31.99±2.04 / 65.82±2.07 28.03±1.96 / 64.53±2.09
Qwen-7B-Inst. 81.45±1.70 / 52.03±2.18 62.95±2.11 / 56.10±2.17 60.32±2.14 / 57.29±2.16 61.61±2.12 / 54.86±2.17
Qwen-14B 63.79±2.10 / 82.84±1.65 49.75±2.18 / 77.08±1.83 45.88±2.18 / 84.52±1.58 47.57±2.18 / 78.27±1.80
Qwen-14B-Inst. 96.53±0.80 / 71.92±1.96 89.78±1.32 / 67.76±2.04 92.66±1.14 / 73.12±1.94 87.85±1.43 / 71.33±1.97
Table 14: Diff-Bias scores (ambiguous / disambiguated) using different retrieval methods from the full-set, with 95% CI half-widths as subscripts.
@@
Bias Type Setting GPT-3.5-turbo Qwen-7B-Inst. Qwen-14B Qwen-14B-Inst. Qwen-72B-Inst.
Gender w/o RAG 26.52±2.81 / 64.30±3.05 90.48±1.87 / 45.88±3.17 72.62±2.84 / 57.30±3.15 96.32±1.20 / 40.84±3.13 93.29±0.96 / 56.90±2.14
stereo-set 13.31±2.16 / 67.08±2.99 37.55±3.08 / 64.09±3.05 42.42±3.15 / 60.29±3.11 76.73±2.69 / 45.78±3.17 71.21±1.66 / 55.35±2.16
full-set 16.77±2.38 / 67.28±2.99 43.40±3.16 / 62.04±3.09 49.46±3.18 / 61.11±3.10 84.52±2.30 / 48.97±3.18 85.39±1.35 / 54.22±2.17
anti-set 12.45±2.10 / 68.42±2.96 35.50±3.05 / 64.30±3.05 44.05±3.16 / 61.52±3.10 75.97±2.72 / 46.09±3.17 69.38±1.60 / 56.28±2.15
Age w/o RAG 7.08±1.15 / 65.66±2.04 83.69±1.55 / 64.55±2.07 49.61±1.75 / 75.98±1.86 87.21±1.39 / 66.41±2.05 77.20±1.70 / 85.79±1.51
stereo-set 13.53±1.35 / 57.37±2.15 43.46±2.13 / 72.66±1.95 41.70±1.83 / 73.93±1.92 76.91±1.79 / 50.83±2.18 66.06±1.95 / 76.52±1.85
full-set 11.24±1.28 / 60.69±2.12 41.60±2.11 / 75.20±1.89 39.89±1.79 / 77.44±1.82 82.03±1.62 / 57.28±2.15 73.54±1.79 / 79.01±1.77
anti-set 10.65±1.27 / 61.19±2.11 41.31±2.08 / 79.50±1.77 36.13±1.71 / 79.74±1.74 75.24±1.82 / 69.53±1.99 66.21±1.94 / 86.97±1.43
Race w/o RAG 61.24±1.93 / 58.70±2.11 98.97±0.34 / 37.37±2.10 73.77±1.30 / 65.30±2.05 99.87±0.16 / 49.23±2.17 99.81±0.14 / 50.06±2.18
stereo-set 50.78±2.17 / 56.51±2.11 64.65±1.96 / 56.06±2.15 55.60±1.54 / 60.42±2.10 94.08±0.70 / 41.47±2.14 92.19±0.99 / 39.91±2.12
full-set 48.63±2.18 / 61.20±2.08 63.42±1.96 / 58.46±2.13 59.51±1.68 / 63.28±2.07 94.40±0.69 / 45.90±2.16 93.69±0.72 / 45.51±2.16
anti-set 49.68±2.17 / 63.87±2.05 69.27±1.89 / 56.12±2.16 61.27±1.60 / 68.23±2.01 96.03±0.59 / 47.53±2.17 97.79±0.44 / 47.53±2.18
Religion w/o RAG 78.05±1.75 / 57.05±2.08 98.71±0.47 / 56.04±2.16 78.66±1.48 / 73.96±1.85 99.43±0.28 / 67.29±2.04 100.00±0.00 / 79.17±1.77
stereo-set 64.01±2.08 / 57.40±2.12 82.00±1.53 / 55.31±2.15 47.74±1.69 / 78.96±1.77 95.37±0.81 / 64.28±2.08 100.00±0.00 / 63.34±2.09
full-set 56.68±2.16 / 57.92±2.11 78.13±1.68 / 58.65±2.14 46.23±1.64 / 80.63±1.71 95.26±0.81 / 68.65±2.01 99.57±0.21 / 67.92±2.02
anti-set 58.94±2.14 / 60.83±2.09 79.96±1.68 / 54.90±2.14 41.81±1.76 / 84.38±1.57 94.40±0.87 / 67.50±2.03 97.42±0.47 / 70.84±1.98
Table 15: Accuracy scores for the CBBQ dataset (Chinese). Scores for ambiguous and disambiguated contexts are separated by ‘/’. For each model and bias type, the maximum accuracy is highlighted in bold.
Bias Type Setting GPT-3.5-turbo Qwen-7B-Inst. Qwen-14B Qwen-14B-Inst. Qwen-72B-Inst.
Gender w/o RAG 30.52±1.44 / 52.68±1.56 77.56±1.31 / 53.66±1.56 42.56±1.55 / 77.89±1.30 82.31±1.20 / 78.35±1.29 90.13±1.24 / 68.64±2.01
stereo-set 23.31±1.32 / 55.93±1.56 48.29±1.57 / 57.54±1.55 28.71±1.42 / 73.90±1.38 61.84±1.52 / 77.86±1.30 76.26±1.77 / 68.92±2.01
full-set 27.40±1.40 / 56.08±1.56 50.08±1.57 / 58.21±1.55 31.06±1.45 / 75.23±1.35 67.15±1.47 / 78.20±1.29 80.73±1.67 / 69.71±2.00
anti-set 24.26±1.34 / 56.29±1.55 45.35±1.56 / 58.00±1.55 25.95±1.37 / 77.28±1.31 61.61±1.52 / 80.52±1.24 75.72±1.78 / 71.14±1.97
Age w/o RAG 9.14±1.18 / 53.46±2.17 59.01±2.07 / 40.17±2.12 16.36±1.34 / 53.58±2.17 48.53±2.08 / 52.85±2.17 75.10±1.79 / 54.32±2.17
stereo-set 7.08±1.08 / 50.00±2.18 29.42±1.93 / 43.20±2.15 14.71±1.29 / 51.38±2.18 44.95±2.08 / 47.80±2.17 58.18±2.04 / 52.67±2.18
full-set 7.99±1.17 / 50.10±2.18 30.98±1.97 / 41.82±2.14 13.70±1.23 / 52.30±2.18 48.53±2.08 / 48.26±2.18 64.62±1.96 / 52.39±2.18
anti-set 7.00±1.06 / 52.57±2.18 33.18±2.01 / 43.29±2.15 12.78±1.16 / 51.75±2.18 46.33±2.06 / 49.91±2.18 66.82±1.94 / 53.68±2.17
Table 16: Accuracy scores for the JBBQ dataset (Japanese). Scores for ambiguous and disambiguated contexts are separated by ‘/’. For each model and bias type, the maximum accuracy is highlighted in bold.
Appendix G Detailed Data for Diff-Bias
This section show the detailed Diff-Bias scores for each separate model for the main part in the paper. Table 17, Table 18 and Figure 9 shows each model that contribute to the average score in Figure 3, Table 4 and Figure 4, respectively.
Bias Type Setting GPT-3.5 Llama3-8B-Inst. Qwen-7B-Inst. Qwen-14B Qwen-14B-Inst. Qwen-72B-Inst.
Gender w/o RAG 5.16±4.36 / -9.33±4.35 5.65±4.36 / 1.59±4.36 10.02±4.34 / -3.67±4.36 3.77±4.36 / -7.34±4.35 -2.38±4.36 / -2.38±4.36 -0.45±0.72 / 1.19±9.97
stereo-set 14.53*±4.32 / 7.14*±4.35 14.68*±4.32 / -0.40±4.37 24.01*±4.24 / 0.50±4.37 13.99*±4.32 / -2.68±4.36 4.61±4.36 / 2.68±4.36 12.21±2.01 / -0.50±10.07
full-set 11.31±4.34 / -0.10*±4.37 6.80±4.36 / -3.97±4.36 15.43±4.31 / -2.08±4.36 0.55±4.37 / -4.66±4.36 -3.08±4.36 / -5.95±4.36 5.80±1.83 / -4.76±9.87
anti-set 4.51±4.36 / -3.97±4.36 0.74±4.37 / -6.85±4.36 10.17±4.34 / -10.12±4.34 -4.51±4.36 / -8.93±4.35 -8.43±4.35 / -12.70*±4.33 0.79±2.35 / -5.56±10.12
Age w/o RAG 41.79±2.94 / 5.92±3.23 31.25±3.07 / 8.32±3.22 30.52±3.08 / 3.42±3.23 38.02±2.99 / 7.34±3.22 18.02±3.18 / 8.59±3.22 13.97±1.59 / 4.94±8.78
stereo-set 32.61±3.05 / 8.97±3.22 27.66±3.10 / 10.71±3.21 35.87±3.02 / 3.15±3.23 38.56±2.98 / 7.01±3.22 18.72±3.17 / 9.35±3.22 20.63±2.06 / 5.93±8.74
full-set 29.67±3.09 / 6.63±3.22 19.67±3.17 / 4.13±3.23 30.52±3.08 / 3.75±3.23 27.53±3.11 / 6.96±3.22 7.50±3.22 / 6.09±3.22 11.63±1.62 / 4.62±8.68
anti-set 17.83±3.18 / 6.30±3.22 8.97±3.22 / 2.77±3.23 20.11±3.16 / 3.53±3.23 6.96±3.22 / 6.79±3.22 2.69±3.23 / 3.26±3.23 8.64±1.87 / 4.79±8.73
Race w/o RAG 10.00±4.50 / 3.40±4.52 6.60±4.51 / 1.06±4.52 1.60±4.52 / 2.02±4.52 6.81±4.51 / 2.13±4.52 0.00±4.52 / -3.30±4.52 0.27±0.59 / 2.13±11.71
stereo-set 24.95*±4.38 / 13.30*±4.48 17.55*±4.45 / 9.26±4.50 12.55*±4.48 / 6.17±4.51 19.95*±4.43 / 8.09±4.51 3.88±4.52 / 3.83±4.52 2.50±1.26 / 2.77±11.62
full-set 16.60±4.46 / 8.83±4.50 12.18±4.49 / 6.91±4.51 7.98±4.51 / 4.89±4.51 13.46±4.48 / 3.83±4.52 0.00±4.52 / 0.64±4.52 -0.05±1.04 / 2.77±11.45
anti-set 6.49±4.51 / 5.43±4.51 7.23±4.51 / 4.15±4.52 4.73±4.52 / 6.11±4.51 4.36±4.52 / 2.23±4.52 -0.43±4.52 / -1.17±4.52 -1.38±1.03 / 0.53±11.66
Religion w/o RAG 8.92±5.64 / 4.33±5.65 18.76±5.56 / 7.17±5.64 5.92±5.65 / 3.50±5.65 12.58±5.61 / 5.00±5.65 8.17±5.64 / 2.83±5.66 5.83±2.02 / 4.33±13.81
stereo-set 14.83±5.60 / 12.50±5.61 17.67±5.57 / 8.50±5.64 16.67±5.58 / 5.83±5.65 22.67±5.51 / 7.17±5.64 10.42±5.63 / 9.33±5.63 8.25±2.43 / 5.00±13.46
full-set 8.00±5.64 / 9.00±5.64 10.17±5.63 / 9.17±5.63 12.83±5.61 / 5.50±5.65 12.58±5.61 / 8.83±5.64 8.42±5.64 / 4.17±5.65 5.75±2.58 / 4.17±13.59
anti-set 2.83±5.66 / 5.17±5.65 11.92±5.62 / 8.83±5.64 6.50±5.65 / 3.67±5.65 8.00±5.64 / 5.00±5.65 7.75±5.64 / 2.00±5.66 4.50±2.72 / 4.17±13.97
Table 17: Diff-Bias scores for the ambiguous and disambiguated contexts (separated by ‘/’) for different bias types and models, with document collections of varying social bias levels used for retrieval. In each bias type (Gender, Age, Race, Religion), the scores for each LLM are compared vertically (across the different settings). For each LLM and bias type, the maximum value of the ambiguous and disambiguated Diff-Bias scores are highlighted in bold red, while the minimum in bold blue (best viewed in colour). 95 % CIs that do not overlap with the corresponding w/o RAG setting are indicated by *.
Bias Type Setting GPT-3.5-turbo Qwen-7B-Inst. Qwen-14B Qwen-14B-Inst. Qwen-72B-Inst.
CBBQ
Gender w/o RAG 18.07±6.26 / 8.64±6.34 7.79±6.35 / 3.91±6.36 9.85±6.33 / 0.00±6.37 3.68±6.36 / 1.44±6.37 5.41±2.07 / 13.08±8.67
stereo-set 35.61±5.95 / 16.26±6.28 46.00±5.65 / 23.05±6.19 32.47±6.02 / 13.17±6.31 21.97±6.21 / 17.49±6.27 22.73±3.89 / 14.08±8.72
full-set 13.74±6.31 / 6.79±6.35 25.43±6.16 / 12.76±6.31 7.25±6.35 / 0.00±6.37 6.39±6.35 / -4.12±6.36 4.22±3.16 / 22.50±8.62
anti-set -6.39±6.35 / 6.38±6.35 -4.33±6.36 / -6.58±6.35 -6.60±6.35 / -10.70±6.33 -9.09±6.34 / -13.58±6.31 -15.26±4.25 / 18.42±8.62
Age w/o RAG 27.07±5.67 / 28.71±5.36 11.13±2.31 / -4.11±5.85 18.26±3.98 / -15.62±5.13 5.96±2.09 / -7.04±5.76 15.97±2.62 / -9.11±4.24
stereo-set 17.82±5.59 / 34.08±5.65 23.15±4.38 / 1.17±5.46 22.46±4.35 / -12.89±5.30 17.34±2.71 / -0.68±6.13 23.29±3.18 / 1.08±5.19
full-set 18.56±5.66 / 26.27±5.72 19.44±4.52 / 0.59±5.29 7.96±4.61 / -17.97±4.98 9.48±2.47 / -10.26±6.02 16.02±2.88 / -2.74±4.98
anti-set 11.04±5.75 / 32.33±5.58 1.56±4.68 / -9.18±4.92 1.96±4.78 / -15.53±4.80 -11.28±2.42 / -17.39±5.51 -4.69±3.45 / -7.72±4.04
Race w/o RAG 4.92±4.32 / 9.52±6.87 -0.13±0.59 / 18.96±6.99 4.23±2.86 / 7.44±6.61 -0.13±0.24 / -1.32±7.21 0.07±0.29 / -11.83±7.18
stereo-set 13.41±4.87 / -10.21±6.98 11.66±4.02 / -2.65±7.16 12.89±4.14 / 15.77±6.75 2.15±1.21 / 5.64±7.13 0.39±1.73 / -2.69±7.07
full-set 7.49±5.04 / -13.05±6.91 8.73±4.12 / -3.63±7.10 9.77±4.32 / 16.22±6.60 1.70±1.18 / 0.78±7.20 0.07±1.26 / -2.94±7.20
anti-set 12.70±4.94 / -10.37±6.84 3.00±3.82 / -4.66±7.19 8.14±4.92 / 11.87±6.43 1.37±1.00 / 2.18±7.21 0.39±0.75 / -2.18±7.22
Religion w/o RAG -2.80±4.14 / -5.63±8.54 1.29±1.02 / 24.17±8.60 6.46±3.71 / -16.67±7.48 0.43±0.42 / 17.92±8.19 0.00±0.00 / -11.67±7.19
stereo-set 5.61±5.44 / 4.38±8.67 14.34±3.38 / 5.21±8.83 24.68±5.83 / -13.75±7.14 4.64±1.69 / 1.88±8.51 0.00±0.00 / -13.34±8.54
full-set -3.88±5.99 / -2.08±8.62 9.16±4.01 / -2.29±8.78 13.04±6.22 / -21.25±6.78 2.16±1.75 / -1.46±8.28 0.43±0.42 / -9.17±8.30
anti-set -6.14±5.81 / -2.50±8.55 -3.88±3.95 / 6.88±8.78 5.39±6.71 / -13.75±6.37 -1.30±1.88 / 1.25±8.37 -2.59±1.01 / -8.34±8.09
JBBQ
Gender w/o RAG 1.51±3.13 / -4.75±3.13 1.53±3.13 / -5.06±3.13 8.77±3.12 / -16.00±3.09 -0.72±3.13 / -20.50±3.07 -3.58±1.34 / -2.82±4.11
stereo-set 12.17±3.11 / 2.15±3.13 13.11±3.11 / -7.21±3.13 17.41±3.09 / -9.36±3.12 11.84±3.11 / -19.22±3.08 1.56±2.10 / -2.97±4.10
full-set 11.50±3.11 / 1.84±3.13 10.35±3.12 / -5.98±3.13 11.58±3.11 / -12.93±3.11 4.22±3.13 / -15.59±3.10 -4.30±1.90 / -4.45±4.07
anti-set 3.25±3.13 / 0.31±3.13 10.53±3.12 / -4.65±3.13 9.56±3.12 / -17.94±3.08 3.73±3.13 / -19.79±3.07 -5.68±2.11 / -3.53±4.02
Age w/o RAG 24.57±7.82 / -6.18±8.37 17.47±5.15 / -1.74±8.25 35.48±7.07 / 3.26±8.38 28.68±5.64 / -2.95±8.38 17.01±3.81 / -1.57±8.37
stereo-set 15.63±7.99 / -3.68±8.40 22.61±6.79 / 1.76±8.33 29.04±7.36 / 4.51±8.39 24.36±5.85 / 3.83±8.39 27.30±4.82 / -4.78±8.39
full-set 13.97±7.97 / -5.35±8.39 16.91±6.83 / -0.38±8.30 23.44±7.55 / 1.14±8.40 14.16±5.86 / 1.41±8.41 18.29±4.63 / -4.94±8.39
anti-set 9.56±8.07 / -7.53±8.38 -0.09±6.86 / -1.58±8.34 4.32±7.84 / -1.04±8.40 5.15±6.08 / -2.39±8.41 10.76±4.64 / -3.56±8.38
Table 18: Diff-Bias scores for the CBBQ dataset (Chinese) and JBBQ dataset (Japanese). Scores for ambiguous and disambiguated contexts are separated by /’. For each model and bias type, the maximum value across settings is highlighted in bold red, and the minimum in bold blue.